You are given the following functions:

 (1) I = 41 - 7r 

 (2) S = 5 + 2r 

 (3) I = S

 a) Find r* and I*

 b) If the there is GHȼ27 million-bond financed increase 

in government expenditure, find:

 i) the new r* and I*. How will AD and Y be affected? 

Explain.

 ii) how much of investment is crowded out? How much 

of consumption is affected? Explain.

 iii) redo ii) using loanable funds diagrams.


1
Expert's answer
2021-05-16T17:42:36-0400

a) I = S, so:

41 - 7r = 5 + 2r,

9r = 36,

r* = 4,

I* = 41 - 7×4 = 13.

b) If the there is GHȼ27 million-bond financed increase in government expenditure, then:

i) r* will increase and I* will increase. AD and Y will increase too.

ii) Some part of investment is crowded as a result of increase in government spending. Consumption will increase.

iii) If we use loanable funds diagrams, then the demand for loanable funds will increase, as a result the interest rate r* and the equilibrium quantity of loanable funds will increase.